The sum of three numbers is 15. If the second is subtracted from 6 times the first, the result is 1 less than the third. If twice the first is added to the second, their sum is 3 more than the third. What are the numbers?
Answer by TimothyLamb(4379)

---
x + y + z = 15
6x - y = z - 1
2x + y = z + 3
---
put the system of linear equations into standard form:
---
x + y + z = 15
6x - y - z = -1
2x + y - z = 3
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 2
y = 6
z = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php